OpenStudy (faiqraees):

tan x = -k Find tan (0.5 pi +x) given that x is obtuse

OpenStudy (faiqraees):

@Photon336 @jhonyy9 @vishweshshrimali5 @ganeshie8 @Directrix @Isaiah.Feynman @IrishBoy123 @inkyvoyd @aaronq @Astrophysics @kropot72 @ParthKohli

Parth (parthkohli):

why not just try the addition formula?

OpenStudy (faiqraees):

That will give infinity in denominator

Parth (parthkohli):

welp, didn't think of that.\[\tan(\pi/2 + x) = -\cot x\]

OpenStudy (faiqraees):

Yeah but that requires me to prove that identity and the question doesnt ask to do it

OpenStudy (faiqraees):

The method basically revolves around using the 4 quadrants and trying to figure out the result

Parth (parthkohli):

\[\frac{\tan \pi/2 + \tan x}{1-\tan \pi/2 \tan x}\]\[= \frac{1+\frac{\tan x}{\tan \pi/2}}{\frac{1}{\tan \pi/2} - \tan x}\]But \(1/\tan \pi/2 =\cot\pi/2 = 0\).

OpenStudy (faiqraees):

I know the identity but the question prohibits me to use it

OpenStudy (faiqraees):

 Its like doing this

Parth (parthkohli):

Yeah, that's how I think about trig. :P

OpenStudy (faiqraees):

Since the tan (p/b) will give k and the quadrant is second so the answer is -k

OpenStudy (faiqraees):

but I dont know how to use the same technique for an obtuse triangle

OpenStudy (faiqraees):

angle*

Parth (parthkohli):

well, you only need to know how to find the sines and cosines really

OpenStudy (faiqraees):

and how would I do that(on a similar note: If I can figure out the sines and cosines they would require the same technique to figure out the tan. So shouldn't I just focus on tan)

Parth (parthkohli):

yeah, but there's no nice way to see tan on a unit circle. ok maybe you can.

OpenStudy (faiqraees):

Any idea how to do that?

Parth (parthkohli):

I don't know. Can you show that \(\sin (\pi/2 + x) = -\cos x\) for an obtuse angle? (Actually any angle.)

OpenStudy (faiqraees):

Actually we haven't been taught the addition formula (although I know it) so I cant use it

Parth (parthkohli):

Ugh, typo. I meant \(\sin (\pi/2 + x) = \cos x\). And no, without the addition formula.

OpenStudy (faiqraees):

Can you think of some way to utilise the quadrant technique this is so far I have gotten